# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ Abstract
Generally, the calculation and memory space required for multi-agent path finding (MAPF) grows exponentially as the number of agents increases. This often results in some MAPF instances being unsolvable under limited computational resources and memory space, thereby limiting the application of MAPF in complex scenarios. Hence, we propose a decomposition approach for MAPF instances, which breaks down instances involving a large number of agents into multiple isolated subproblems involving fewer agents. Moreover, we present a framework to enable general MAPF algorithms to solve each subproblem independently and merge their solutions into one conflict-free final solution, without compromising on solvability. Unlike existing works that propose isolated methods aimed at reducing the time cost of MAPF, our method is applicable to all MAPF methods. In our results, we apply decomposition to multiple state-of-the-art MAPF methods using a classic MAPF benchmark [^1] https://movingai.com/benchmarks/mapf.html. The decomposition of MAPF instances is completed on average within 1s, and its application to seven MAPF methods reduces the memory usage and time cost significantly, particularly for serial methods. To facilitate further research within the community, we have made the source code of the proposed algorithm publicly available [^2] https://github.com/JoeYao-bit/LayeredMAPF.
Index Terms:
Multi-Agent Path Finding, graph theory.
I Introduction
Multi-agent pathfinding (MAPF), as its name suggests, computes a set of collision-free paths for multiple agents from their respective starting locations to target locations.
@@ I Introduction
Motivated by the phenomenon that the cost of solving MAPF instances grows exponentially as the number of agents increases [li2022mapf], we propose a novel approach to reduce the cost of MAPF methods by decomposing a MAPF instance into multiple smaller subproblems. These subproblems are solved independently, while considering the solutions of other subproblems as dynamic obstacles.
This idea bears resemblance to Priority-Based Search (PBS) [ma2019searching], which assigns a unique priority to each agent and solves agents separately in decreasing priority order. PBS can be viewed as decomposing a MAPF instance into subproblems, each involving only one agent. While PBS is efficient, it lacks a guarantee of completeness. In contrast, our approach aims to decompose MAPF instances without compromising their solvability.
We formulate the decomposition of MAPF instances as a progressive optimization problem. Initially, we decompose the MAPF instance into multiple subproblems without restricting the order of solving, and then further split them into smaller subproblems with a limited solving order. To ensure solvability and minimize the size of each decomposition, we evaluate the solvability of each step of decomposition, allowing only those decompositions that do not compromise solvability. As a result, we demonstrate the performance of our method across various maps and illustrate its improvement on multiple cutting-edge MAPF methods.
Figure 1: These figures illustrate the complete process of decomposing a MAPF instance, which includes decomposing agents into initial clusters (a type of subproblem), further decomposing initial clusters into smaller clusters, and decomposing clusters into levels (another type of subproblem). For more details, please refer to Section IV.
@@
1. We introduce a novel problem of decomposing a MAPF instance into solvable subproblems to reduce the solving cost. Additionally, we propose a novel method that decomposes a MAPF instance into multiple subproblems without compromising solvability. An overview of our method is depicted in Fig. 1.
@@
The remainder of this article is organized as follows. In Section II, we review related problems and methods aimed at reducing the solving cost of MAPF instances. We then define basic terminology in Section III and introduce our method in Section IV. Our test results regarding the performance of decomposition under various maps are presented in Section V, followed by an examination of how decomposition benefits multiple cutting-edge MAPF methods in Section VI. Finally, Section VII concludes this article.
II Related works
@@
1, move one agent from its start to target while keeping other agents static and attempting to avoid conflicts with the paths of other agents. This approach is adopted by CBS-based methods, HCA*, LNS, and PBS.
2, move all agents simultaneously by avoiding conflicts with each other at each time step. This approach is employed by LaCAM, PIBT, and Push and Swap.
We refer to the first approach as serial MAPF methods and the second approach as parallel MAPF methods. This distinction is crucial in understanding how the decomposition of a MAPF instance applies to them. Further details about this difference can be found in Section IV-D.
II-A Conflict based search
Conflict-Based Search (CBS) [sharon2015conflict] is a two-level complete and optimal multi-agent pathfinding algorithm. At the high level, CBS employs a conflict tree in which nodes represent conflicts between paths of two agents. By splitting a conflict, CBS obtains constraints for the related two agents, respectively. The low level utilizes heuristic single-agent searches to find new paths that satisfy all constraints imposed by the high-level conflict tree node. Therefore, CBS operates as a serial MAPF method. If CBS finds a new path, it checks for conflicts between the new path and previous paths. If no new conflicts are found, CBS exits with a conflict-free solution; otherwise, it inserts the new conflict as nodes into the conflict tree and repeats the process of splitting conflicts and searching for new paths.
@@ II-A Conflict based search
II-A1 Trade off solution quality for efficency
@@
Bounded CBS (BCBS) [hatem2013bounded] incorporates focal search into the low-level search of CBS, enabling the low-level search to consider avoiding conflicts with other agents’ paths and generate bounded suboptimal solutions. Enhanced CBS (ECBS) [hatem2013bounded] utilizes the same low-level search as BCBS and applies focal search to the high-level search of CBS, aiming to minimize the number of nodes from the current conflict tree (CT) node to the CT node representing the conflict-free solution. Similar to BCBS, ECBS also produces bounded suboptimal solutions.
@@
Explicit Estimation CBS (EECBS) [li2021eecbs] employs online learning to estimate the cost of the solution beneath each high-level node. It utilizes Explicit Estimation Search (EES) [thayer2011bounded] to select high-level nodes for expansion, mitigating two drawbacks of ECBS: (1) the cost of the solution beneath a CT node $N$ might exceed the cost of $N$ and thus could surpass the suboptimality bound; (2) the lower bound of ECBS seldom increases, leading to difficulties in finding a solution within a reasonable time if the optimal sum of costs is not within the initial suboptimality bound.
II-A2 Bypassing conflicts
@@
Bypassing Conflicts [boyarski2015don] is a conflict resolution technique that alters the paths of agents involved in a chosen conflict instead of splitting a conflict tree (CT) node and searching for paths to avoid related constraints. When expanding a CT node $N$ and generating child CT nodes, if the cost of a child CT node equals the cost of $N$ and the number of conflicts of the paths in the child node is smaller than the number of conflicts of $N$, then the child node is used to replace $N$, and all generated child CT nodes are discarded. Otherwise, $N$ is split as before. It has been demonstrated that bypassing conflicts often results in smaller CTs and reduces the runtime of CBS.
II-A3 Prioritizing conflicts
@@
Prioritizing Conflicts [boyarski2015don] is a conflict-selection technique. A conflict is considered cardinal if, when CBS uses the conflict to split a conflict tree (CT) node $N$, the costs of both child CT nodes are larger than the cost of $N$. It is semi-cardinal if the cost of one child CT node is larger than the cost of $N$, and the cost of the other node is not. It is non-cardinal if the costs of both child CT nodes are equal to the cost of $N$. CBS can significantly enhance its efficiency by resolving cardinal conflicts first, then semi-cardinal conflicts, and finally non-cardinal conflicts. This prioritization is effective because generating CT nodes with larger costs first typically improves the lower bound of the CT (i.e., the minimum cost of the CT nodes in the open list) faster, thereby producing a smaller conflict tree and accelerating CBS.
II-A4 Symmetry reasoning
S
@@ II-A4 Symmetry reasoning
Symmetry Reasoning [li2019symmetry, li2020new] is a technique aimed at avoiding the repeated resolution of conflicts between the same pair of agents due to symmetric paths and conflicts. It efficiently identifies each symmetry and resolves it through a single splitting action with specialized constraints. This process results in a smaller conflict tree and reduces the time cost of CBS.
II-A5 Weighted Dependency Graph (WDG)
@@
The Weighted Dependency Graph (WDG) heuristic [li2019improved] is an admissible heuristic employed in the high-level search of CBS. It operates by constructing a weighted dependency graph for each conflict tree (CT) node $N$. In these graphs, vertices represent agents, and edges denote dependency between agents. This dependency is defined as follows: the minimum sum of costs of their conflict-free paths satisfying $N$’s constraints (computed via solving a 2-agent MAPF instance using CBS) is greater than the sum of costs of their paths in $N$’s paths (the shortest paths satisfying $N$’s constraints but not necessarily conflict-free). The edge weights signify the difference between the minimum sum of costs of conflict-free paths satisfying $N$’s constraints and the sum of costs of their paths in $N$’s paths.
@@
The value of the edge-weighted minimum vertex cover of the graph serves as an admissible heuristic for the high-level search of CBS. Despite the runtime overhead associated with building the weighted dependency graphs and determining their edge-weighted minimum vertex cover, incorporating the WDG heuristic often leads to a reduction in conflict tree size and a decrease in the runtime of CBS.
II-B Large neighborhood search
At one end of the spectrum are (bounded-sub) optimal algorithms capable of finding high-quality solutions for small problems. At the opposite end are unbounded-suboptimal algorithms, which can handle large problems but typically produce low-quality solutions. Li et al. proposed MAPF-LNS [li2021anytime], presenting a third approach that combines the strengths of both.
@@ II-B Large neighborhood search
MAPF-LNS starts with a given solution and deletes a portion of it, referred to as a neighborhood, while treating the remaining part as fixed. This simplifies the original problem, making it easier to solve. If the newly found solution is superior to the current one, it replaces the existing solution. Consequently, MAPF-LNS is both complete and capable of iteratively enhancing solution quality towards near-optimality. During each iteration, LNS selects a subset of agents and updates their paths without modifying those of other agents. Thus, MAPF-LNS operates as a serial MAPF method. MAPF-LNS can employ any desired approach to solve the simplified problem, provided it can account for the fixed information.
@@
Building on this work, Li et al. introduced MAPF-LNS2 [li2022mapf], which efficiently finds a solution (rather than improving a given solution) for a MAPF instance. Initially, MAPF-LNS2 invokes a MAPF algorithm to solve the instance and acquire a (partial or complete) plan. For agents lacking a path, MAPF-LNS2 devises a path that minimizes collisions with existing paths.
II-C Priority based search
@@
Priority-Based Search (PBS) [ma2019searching] is an incomplete, suboptimal algorithm designed for prioritized planning. It employs a depth-first search at the high level to dynamically construct a priority ordering, thereby forming a priority tree (PT). When confronted with a collision, PBS greedily determines which agent should be assigned a higher priority. It efficiently backtracks and explores alternative branches only if no solution is found in the current branch. Consequently, PBS incrementally constructs a single partial priority ordering until all collisions are resolved.
@@
PBS ranks among the most efficient methods for solving MAPF. However, prioritized planning with an arbitrary priority ordering does not guarantee completeness or optimality in general [ma2019searching].
II-C1 Greedy PBS
Given that PBS becomes less effective for MAPF instances with high densities of agents and obstacles, [chan2023greedy] introduced Greedy PBS (GPBS). GPBS utilizes greedy strategies to enhance PBS by minimizing the number of collisions between agents. In essence, GPBS employs the numbers of conflicts and conflicting pairs as heuristics to guide the search on its low and high levels, respectively.
PBS can be regarded as a special decomposition of MAPF instances, where each subproblem involves only one agent. Unlike PBS, our decomposition does not necessitate that every subproblem contain only one agent, while still preserving completeness. Additionally, our decomposition serves as an auxiliary component rather than an isolated method, making it applicable to a wide range of MAPF algorithms.
II-D PIBT and LaCAM
II-D1 Priority Inheritance with Backtracking (PIBT)
PIBT [okumura2019priority] is an incomplete and suboptimal MAPF method that assigns a unique priority to each agent at every timestep. This prioritization ensures that all movements are ordered, and all agents act within a single timestep according to their priority, making it a parallel MAPF method. Priority inheritance is utilized to effectively handle priority inversion in path adjustment within a small time window, and a backtracking protocol prevents agents from becoming stuck.
@@ II-D1 Priority Inheritance with Backtracking (PIBT)
In PIBT, for each timestep, each agent must evaluate distances from surrounding nodes to its goal. While this operation could be implemented by calling A* on demand, it may also become a bottleneck. To address this issue, the authors of PIBT proposed PIBT+ [okumura2022priority], which saves computation time by preparing distance tables from each agent’s goal.
II-D2 Lazy constraints addition for MAPF (LaCAM)
LaCAM [okumura2023lacam] is a complete and suboptimal two-level MAPF method. At the high level, it explores a sequence of configurations, where each search node corresponds to one configuration. At the low level, it searches for constraints specifying which agents go where in the next configuration, making it a parallel MAPF method.

### v2

# LayeredMAPF: a decomposition of MAPF instance to reduce solving costs

@@ Abstract
Multi-agent pathfinding (MAPF) holds significant utility within autonomous systems, however, the calculation and memory space required for multi-agent path finding (MAPF) grows exponentially as the number of agents increases. This often results in some MAPF instances being unsolvable under limited computational resources and memory space, thereby limiting the application of MAPF in complex scenarios. Hence, we propose a decomposition approach for MAPF instances, which breaks down instances involving a large number of agents into multiple isolated subproblems involving fewer agents. Moreover, we present a framework to enable general MAPF algorithms to solve each subproblem independently and merge their solutions into one conflict-free final solution, and avoid loss of solvability as much as possible. Unlike existing works that propose isolated methods aimed at reducing the time cost of MAPF, our method is applicable to all MAPF methods. In our results, we apply decomposition to multiple state-of-the-art MAPF methods using a classic MAPF benchmark [^1] https://movingai.com/benchmarks/mapf.html. The decomposition of MAPF instances is completed on average within 1s, and its application to seven MAPF methods reduces the memory usage or time cost significantly, particularly for serial methods. Based on massive experiments, we speculate the possibilty about loss of solvability caused by our method is $<$ 1%. To facilitate further research within the community, we have made the source code of the proposed algorithm publicly available [^2] https://github.com/JoeYao-bit/LayeredMAPF/tree/minimize_dependence.
keywords:
Multi-Agent Path Finding (MAPF), graph theory
1 Introduction
Multi-agent pathfinding (MAPF), as its name suggests, computes a set of collision-free paths for multiple agents from their respective starting locations to target locations. MAPF is widely utilized in autonomous systems, such as automated warehouses [1] and UAV traffic management [2].
@@ 1 Introduction
Motivated by the phenomenon that the cost of solving MAPF instances grows exponentially as the number of agents increases [3], we propose a novel approach to reduce the cost of MAPF methods by decomposing a MAPF instance into multiple smaller subproblems. These subproblems are solved independently, while considering the solutions of other subproblems as dynamic obstacles.
This idea bears resemblance to Priority-Based Search (PBS) [4], which assigns a unique priority to each agent and solves agents separately in decreasing priority order. PBS can be viewed as decomposing a MAPF instance into subproblems, each involving only one agent. While PBS is efficient, it lacks a guarantee of completeness. Compring to PBS, our method try to reduce possibility of loss of solvability by make each subproblem contain more than one agent. We formulate the decomposition of MAPF instances as a progressive optimization problem. Initially, we decompose the MAPF instance into multiple subproblems without restricting the order of solving, and then further split them into smaller subproblems with a limited solving order. To minimize loss of solvability and minimize the size of each subproblem, we evaluate the solvability of each step of decomposition, allowing only decompositions that pass a solvability check. As a result, we demonstrate the performance of our method across various maps and illustrate its improvement on multiple cutting-edge MAPF methods.
Figure 1: These figures illustrate the complete process of decomposing a MAPF instance, which includes decomposing agents into initial clusters (a type of subproblem), further decomposing initial clusters into smaller clusters, and decomposing clusters into levels (another type of subproblem). For more details, please refer to Section 4.
@@
1. We propose a novel method that decomposes a MAPF instance into multiple subproblems with minimizing loss of solvability. An overview of our method is depicted in Fig. 1.
@@
The remainder of this article is organized as follows. In Section 2, we review related problems and methods aimed at reducing the solving cost of MAPF instances. We then define basic terminology in Section 3 and introduce our method in Section 4. Our test results regarding the performance of decomposition under various maps are presented in Section 5, followed by an examination of how decomposition benefits multiple cutting-edge MAPF methods in Section 6. Finally, Section 7 concludes this article.
2 Related works
@@
1, plan one agent’s path from its start to target while keeping other agents static and attempting to avoid conflicts with the paths of other agents. This approach is adopted by CBS-based methods, HCA*, LNS, and PBS.
2, plan for all agents simultaneously by avoiding conflicts with each other at each time step. This approach is employed by LaCAM, PIBT, and Push and Swap.
We refer to the first approach as serial MAPF methods and the second approach as parallel MAPF methods. This distinction is crucial in understanding how the decomposition of a MAPF instance applies to them. Further details about this difference can be found in Section 4.4.
2.1 Conflict based search
Confict-Based Search (CBS) [5] is a two-level complete and optimal multi-agent pathfinding algorithm. At the high level, CBS employs a constraint tree in which nodes represent conflicts between paths of two agents. By splitting a conflict, CBS obtains constraints for the related two agents, respectively. The low level utilizes heuristic single-agent searches to find new paths that satisfy all constraints imposed by the high-level constraint tree node. Therefore, CBS operates as a serial MAPF method. If CBS finds a new path, it checks for conflicts between the new path and previous paths. If no new conflicts are found, CBS exits with a conflict-free solution; otherwise, it inserts the new constraints generated from a conflict as nodes into the conflict tree and repeats the process of splitting conflicts and searching for new paths. If no new conflicts are found, CBS exits with a conflict-free solution. If conflicts exist in the current high-level node, the algorithm first selects one conflict, then generates a pair of constraints to resolve the conflict. Finally, it generates two child nodes, each adding a new constraint (together with existing constraints inherited from the current node). Then in each child node, paths are then replanned to satisfy new constraints and conflicts are updated.
@@ 2.1 Conflict based search
$\bullet$ Trade off solution quality for efficency
@@
Bounded CBS (BCBS) [6] incorporates focal search into the low-level search of CBS, enabling the low-level search to consider avoiding conflicts with other agents’ paths and generate bounded suboptimal solutions. Enhanced CBS (ECBS) [6] utilizes the same low-level search as BCBS and applies focal search to the high-level search of CBS, aiming to minimize the number of nodes from the current constraint tree (CT) node to the CT node representing the conflict-free solution. Similar to BCBS, ECBS also produces bounded suboptimal solutions.
@@
Explicit Estimation CBS (EECBS) [7] employs online learning to estimate the cost of the solution beneath each high-level node. It utilizes Explicit Estimation Search (EES) [8] to select high-level nodes for expansion, mitigating two drawbacks of ECBS: (1) the cost of the solution beneath a CT node $N$ might exceed the cost of $N$ and thus could surpass the suboptimality bound; (2) the lower bound of ECBS seldom increases, leading to difficulties in finding a solution within a reasonable time if the optimal sum of costs is not within the initial suboptimality bound.
$\bullet$ Bypassing conflicts
@@
Bypassing Conflicts [9] is a conflict resolution technique that alters the paths of agents involved in a chosen conflict instead of splitting a constraint tree (CT) node and searching for paths to avoid related constraints. When expanding a CT node $N$ and generating child CT nodes, if the cost of a child CT node equals the cost of $N$ and the number of conflicts of the paths in the child node is smaller than the number of conflicts of $N$, then the child node is used to replace $N$, and all generated child CT nodes are discarded. Otherwise, $N$ is split as before. It has been demonstrated that bypassing conflicts often results in smaller CTs and reduces the runtime of CBS.
$\bullet$ Prioritizing conflicts
@@
Prioritizing Conflicts [9] is a conflict-selection technique. A conflict is considered cardinal if, when CBS uses the conflict to split a constraint tree (CT) node $N$, the costs of both child CT nodes are larger than the cost of $N$. It is semi-cardinal if the cost of one child CT node is larger than the cost of $N$, and the cost of the other node is not. It is non-cardinal if the costs of both child CT nodes are equal to the cost of $N$. CBS can significantly enhance its efficiency by resolving cardinal conflicts first, then semi-cardinal conflicts, and finally non-cardinal conflicts. This prioritization is effective because generating CT nodes with larger costs first typically improves the lower bound of the CT (i.e., the minimum cost of the CT nodes in the open list) faster, thereby producing a smaller constraint tree and accelerating CBS.
$\bullet$ Symmetry reasoning
@@
Symmetry Reasoning [10, 11] is a technique aimed at avoiding the repeated resolution of conflicts between the same pair of agents due to symmetric paths and conflicts. It efficiently identifies each symmetry and resolves it through a single splitting action with specialized constraints. This process results in a smaller constraint tree and reduces the time cost of CBS.
$\bullet$ Weighted Dependency Graph (WDG)
@@
The Weighted Dependency Graph (WDG) heuristic [12] is an admissible heuristic employed in the high-level search of CBS. It operates by constructing a weighted dependency graph for each constraint tree (CT) node $N$. In these graphs, vertices represent agents, and edges denote dependency between agents. This dependency is defined as follows: the minimum sum of costs of their conflict-free paths satisfying $N$’s constraints (computed via solving a 2-agent MAPF instance using CBS) is greater than the sum of costs of their paths in $N$’s paths (the shortest paths satisfying $N$’s constraints but not necessarily conflict-free). The edge weights signify the difference between the minimum sum of costs of conflict-free paths satisfying $N$’s constraints and the sum of costs of their paths in $N$’s paths.
@@
The value of the edge-weighted minimum vertex cover of the graph serves as an admissible heuristic for the high-level search of CBS. Despite the runtime overhead associated with building the weighted dependency graphs and determining their edge-weighted minimum vertex cover, incorporating the WDG heuristic often leads to a reduction in constraint tree size and a decrease in the runtime of CBS.
2.2 Large neighborhood search
Large Neighborhood Search (LNS) is a classic algorithm for finding good solutions to challenging discrete optimization problems. Li et al. [13] proposed the first (to our best knowledge) LNS approaches for MAPF, MAPF-LNS.
@@ 2.2 Large neighborhood search
MAPF-LNS starts with a given solution and deletes a portion of it, referred to as a neighborhood, while treating the remaining part as fixed. This simplifies the original problem, making it easier to solve. If the newly found solution is superior to the current one, it replaces the existing solution. Consequently, MAPF-LNS is capable of iteratively enhancing solution quality towards near-optimality. Its completeness depends on the algorithm used to generate initial paths. During each iteration, LNS selects a subset of agents and updates their paths without modifying those of other agents. Thus, MAPF-LNS operates as a serial MAPF method. MAPF-LNS can employ any desired approach to solve the simplified problem, provided it can account for the fixed information.
@@
Building on this work, Li et al. introduced MAPF-LNS2 [3], which efficiently finds a solution (rather than improving a given solution) for a MAPF instance. Initially, MAPF-LNS2 invokes a MAPF algorithm to solve the instance and acquire a (partial or complete) plan. For agents lacking a path, MAPF-LNS2 devises a path that minimizes collisions with existing paths.
2.3 Priority based search
@@
Priority-Based Search (PBS) [4] is an incomplete, suboptimal algorithm designed for prioritized planning. It employs a depth-first search at the high level to dynamically construct a priority ordering, thereby forming a priority tree (PT). When confronted with a collision, PBS greedily determines which agent should be assigned a higher priority. It efficiently backtracks and explores alternative branches only if no solution is found in the current branch. Consequently, PBS incrementally constructs a single partial priority ordering until all collisions are resolved.
@@
PBS ranks among the most efficient methods for solving MAPF. However, prioritized planning with an arbitrary priority ordering does not guarantee completeness or optimality in general [4].
$\bullet$ Greedy PBS
Given that PBS becomes less effective for MAPF instances with high densities of agents and obstacles, [14] introduced Greedy PBS (GPBS). GPBS utilizes greedy strategies to enhance PBS by minimizing the number of collisions between agents. In essence, GPBS employs the numbers of conflicts and conflicting pairs as heuristics to guide the search on its low and high levels, respectively.
PBS can be regarded as a special decomposition of MAPF instances, where each subproblem involves only one agent. Unlike PBS, our decomposition does not necessitate that every subproblem contain only one agent, so our decomposition has less loss of solvability. Additionally, our decomposition serves as an auxiliary component rather than an isolated method, making it applicable to a wide range of MAPF algorithms.
2.4 PIBT and LaCAM
2.4.1 Priority Inheritance with Backtracking (PIBT)
PIBT [15] is an incomplete and suboptimal MAPF method that assigns a unique priority to each agent at every timestep. This prioritization ensures that all movements are ordered, and all agents act within a single timestep according to their priority, making it a parallel MAPF method. Priority inheritance is utilized to effectively handle priority inversion in path adjustment within a small time window, and a backtracking protocol prevents agents from becoming stuck.
@@ 2.4.1 Priority Inheritance with Backtracking (PIBT)
In PIBT, for each timestep, each agent must evaluate distances from surrounding nodes to its goal. While this operation could be implemented by calling A* on demand, it may also become a bottleneck. To address this issue, the authors of PIBT proposed PIBT+ [16], which saves computation time by preparing distance tables from each agent’s goal.
2.4.2 Lazy constraints addition for MAPF (LaCAM)
LaCAM [17] is a complete and suboptimal two-level MAPF method. At the high level, it explores a sequence of configurations, where each search node corresponds to one configuration. At the low level, it searches for constraints specifying which agents go where in the next configuration, making it a parallel MAPF method.